# Using diffusion model as constraint: Empower Image Restoration Network Training with Diffusion Model

### v1

@@ 5 Experiments
FFA-Net[46] 14.39 0.4524 RESCAN[35] 26.36 0.7860 SwinIR[36] 17.81 0.7230
AECR-Net[64] 15.80 0.4660 PreNet[48] 26.77 0.8580 AirNet[34] 18.18 0.7350
@@ 5.1 Experiment Setup
Datasets. We train and evaluate our models on both synthetic and real-world image restoration datasets, including low light enhancement, image deraining and image dehazing. For real-world challenging scenes, we adopt LOL dataset[61] for low light enhancement and Dense-Haze dataset[1] for image dehazing. For image deraining, we adopt Rain13K dataset[6] for training and Rain100H dataset[65] for testing.
Finally, we use CUB dataset[56] for image classification task, and the degraded CUB dataset is obtained through synthetic methods from [25] to simulate degraded conditions, such as fog, rain, and low-light scenarios.
@@
Figure 7: Comparison of visual results on Dense-Haze dataset. (e): FSDGN; (f): Taylorformer. Please zoom in for best view.
@@
Evaluation Metrics. We evaluate our method on two different metrics: PSNR, and SSIM [59], which well-known image quality assessment indicators.
@@
We also show the visual comparison with other method on real degradation images sampled from testing sets of Dense-Haze, Rain100H and LOL datasets. The visual results are presented in  Fig. 6 and Fig. 7. The complex degradation distribution makes these datasets extremely challenging, and existing methods on these datasets usually suffer from unnaturalness problems. As shown in  Fig. 6 and Fig. 7, these baselines produce relatively pleasing results, however, artifacts and blurs still emerge. In contrast, after being empowered by the DiffLoss, they produce results more visually pleasing than the baseline results. For better visualization, we denote the obvious region with the red rectangular box.
@@
Improvement on Image Classification. We have demonstrated in Fig. 2 that h-space possesses semantic attributes. Therefore, we utilize h-space as a loss to preserve the semantic information for low level tasks. We synthesized degraded images with low-light, haze, and rain using the origin CUB dataset. We train the model using common low-level datasets as mentioned before with and without DiffLoss, and use the trained model to restore the degraded CUB dataset. We choose to use Taylorformer, EfDeRain and IAT for different degradations. Finally, we employ a pre-trained VGG16[54] and Resnet50[24] networks to evaluate the accuracy of classification. From the Tab. 2, it is evident that with the assistance of our method, the model can preserve more semantic information after the restoration process.
Figure 8: Graph of PSNR with different weight of DiffLoss during training process on Dense-Haze [1] dataset with MSBDN [14] as baseline. We train the baseline from scratch and choose the best performance in the first 40K iterations.
@@ 5.3 Ablation study
In this section, we perform several ablation studies to analyze the effectiveness of the proposed DiffLoss on Dense-Haze dataset. The studies include the following ablated models: (a) With constraints on $\hat{x}_{0}$ only. (b) With constraints on $\hat{\hat{x}}_{0}$ only. (c) With constraints on $\hat{\varepsilon}$ and $\hat{x}_{t-1}$. (d) Ours (final setting). These models are trained using the same training setting as our method. The performance of these models is summarized in Table 3. Obviously, every design component in DiffLoss can elevate the performance, and constraining on $\hat{\varepsilon}$ and $\hat{x}_{t-1}$ performs similarly to constraining on $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$.
We also experiment with different loss value to get the optimal one. As shown in Fig. 8, we experiment with different weight of DiffLoss and plot the PSNR-Loss weight curve. Similar to perceptual and adversarial loss, DiffLoss gets the optimal performance with loss weight $\gamma\in[0.0005,0.005]$. In our method, we choose the loss weight of DiffLoss to be 0.001. This experiment is conducted with $\lambda=0$. The experiment of optimal choice for $\lambda$ and more comparison presented in the supplementary material.
@@ 6 Conclusion
In this paper, we propose a new perspective from the correlation of degraded and natural image distribution that achieves effective image restoration. To achieve this, inspired by the powerful capability of the diffusion model for natural image sampling and generation, we embed the pre-trained diffusion model into restoration network as a auxiliary training mechanism to empower the learning capability and semantic attributes of neural networks for effective naturalness image restoration.
By equipping existing restoration networks with the DiffLoss in the training stage, we can substantially elevate their performance and yield more natural and recognition-aware restored images without involving additional computations in the inference stage.

### v2

@@ 4 Experiments
FFA-Net[46] 14.39 0.4524 RESCAN[37] 26.36 0.7860 RetiNexNet[60] 16.77 0.5620
AECR-Net[64] 15.80 0.4660 PreNet[48] 26.77 0.8580 DRBN[65] 19.55 0.7460
@@ 4.1 Experiment Setup
Datasets. We train and evaluate our models on both synthetic and real-world image restoration datasets, including low light enhancement, image deraining and image dehazing. For real-world challenging scenes, we adopt LOL dataset[61] for low light enhancement and Dense-Haze dataset[1] for image dehazing. For image deraining, we adopt Rain13K dataset[6] for training and Rain100H dataset[66] for testing. Finally, we use CUB dataset[56] for image classification task, and the degraded CUB dataset is obtained through synthetic methods from [26] to simulate degraded conditions, such as fog, rain, and low-light scenarios.
@@
Figure 6: Comparison of visual results on Dense-Haze dataset. We also show the visual comparison between different loss functions and our method.
@@
Evaluation Metrics. We evaluate our method on two different metrics: PSNR, SSIM [59], LPIPS[72] and FID[70] which are well-known image quality assessment indicators.
@@
We also show the visual comparison with other method on real degradation images sampled from testing sets of Dense-Haze, Rain100H and LOL datasets. The visual results are presented in  Fig. 5 and Fig. 6. The complex degradation distribution makes these datasets extremely challenging, and existing methods on these datasets usually suffer from unnaturalness problems. As shown in  Fig. 5 and Fig. 6, these baselines produce relatively pleasing results, however, artifacts and blurs still emerge. In contrast, after being empowered by the DiffLoss, they produce results more visually pleasing than the baseline results. We also present a comparison with different loss functions in Fig. 6, which is described in the following sections. For better visualization, we denote the obvious region with the red rectangular box.
@@
Table 3: The performance comparison of different loss functions on Dense-Haze [1] dataset with MSBDN [14] as baseline. We train the baseline from scratch and choose the best performance in the first 40K iterations.
[…]
L1 Loss 0.4948 15.228 0.4974
Perceptual Loss 0.4921 15.609 0.5011
Adversarial Loss 0.4848 15.593 0.4981
DiffLoss 0.4731 15.682 0.5088
Improvement on Image Classification. We have demonstrated in Fig. 2 that h-space possesses semantic attributes. Therefore, we utilize h-space as a loss to preserve the semantic information for low level tasks. We synthesized degraded images with low-light, haze, and rain using the origin CUB dataset. We train the model using common low-level datasets as mentioned before with and without DiffLoss, and use the trained model to restore the degraded CUB dataset. We choose to use Taylorformer, EfDeRain and IAT for different degradations. Finally, we employ a pre-trained VGG16[54] and Resnet50[25] networks to evaluate the accuracy of classification. From the Tab. 2, it is evident that with the assistance of our method, the model can preserve more semantic information after the restoration process. Please note that our focus is not on comparing with other methods, but rather on improving existing approaches.
Comparison with other loss functions. Previous loss [12, 17] that assist restoration networks have the following drawbacks: (1) L1 or L2 loss works in pixel space, which may produce images deviated from natural distribution. (2) The VGG16 used in perceptual loss is pre-trained for high-level task, instead of low-level image restoration task. (3) Adversarial loss treats restoration network as generator and inserts an additional discriminator network and needs to train a discriminator for every restoration dataset, which is troublesome and time-consuming. The wide distribution and mode convergence property enables diffusion model to be a powerful and general image prior, and fits for both general and specific low-level image restoration tasks. Besides, h-space in diffusion model is also found to reflect the semantics of images. As shown in Fig. 6 and Tab. 3, our DiffLoss demonstrates excellent performance compared to other approaches.
By pulling the intermediate sampling stages and h-space closer to that of clear images and leveraging the distribution sampling property of diffusion model, the restored results can be optimized to be more natural and recognition-aware, which is difficult to achieve for the previous methods.
Figure 7: Graph of PSNR and Top-1V(%) with different weight of DiffLoss during training process on Dense-Haze [1] dataset with MSBDN [14] as baseline. We train the baseline from scratch and choose the best performance in the first 40K iterations.
@@ 4.3 Ablation study
In this section, we perform several ablation studies to analyze the effectiveness of the proposed DiffLoss on Dense-Haze dataset. The studies include the following ablated models: (a) With constraints on $\hat{x}_{0}$ only. (b) With constraints on $\hat{x}_{t-1}$ only. (c) Ours (final setting). These models are trained using the same training setting as our method. The performance of these models is summarized in Table 4. Obviously, every design component in DiffLoss can elevate the performance.
We also experiment with different loss value to get the optimal one. As shown in Fig. 7, we experiment with different weight of DiffLoss and plot the PSNR-Loss weight curve. DiffLoss gets the optimal performance with loss weight $\gamma\in[0.0005,0.005]$. In our method, we choose the loss weight of DiffLoss to be 0.001. This experiment is conducted with $\lambda=0$. For the optimal choice of $\lambda$, we find that when $\lambda=0.01$, it achieves the best performance. Note that in the experiments for selecting $\lambda$, we set $\gamma=0.001$.
@@ 5 Conclusion
In this paper, we propose a new perspective from the correlation of degraded and natural image distribution that achieves effective image restoration. To achieve this, inspired by the powerful capability of the diffusion model for natural image sampling and generation, we embed the pre-trained diffusion model into restoration network as a auxiliary training mechanism to empower the learning capability and semantic attributes of neural networks for effective naturalness image restoration. By equipping existing restoration networks with the DiffLoss in the training stage, we can substantially elevate their performance and yield more natural and semantic-aware restored images without involving additional computations in the inference stage.